Prove weather following statements are true or false.

  1. 7n3+n2 ≠ Ω(n4)
  2. 9n + 3 = O(n2)
  3. n*lgn + 8n2 = Ω(n3)
  4. 2n + n3 + 1 = Ө(n3)
1
Expert's answer
2021-11-04T11:44:39-0400

"f(n)=O(g(n))" if there exists a positive real number M and a real number n0 such that

"|f(n)|\le Mg(n)" for "n\ge n_0"

"f(n)=\Omega(g(n))" if there exists a positive real number M and a real number n0 such that

"|f(n)|\ge Mg(n)" for "n\ge n_0"

"f(n)=\theta(g(n))" if there exists a positive real numbers M1 and M2 and a real number n0 such that

"M_1g(n)\le|f(n)|\le M_2g(n)" for "n\ge n_0"


1.

"7n^3+n^2\le 8n^3"

"8n^4\ge 8n^3\ge7n^3+n^2"


2.

"9n+3\le 12n"

"12n^2\ge 12n \ge 9n+3"


3.

"n^3\le nlgn+8n^2"


4.

"n^3\le 2n+n^3+1\le 4n^3"